OpenStudy (anonymous):

FREE MEDAL FOR HELP Which statement is true about the product square root of 2(5 + square root of 8)?

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (anonymous):

\[\sqrt{2(5+\sqrt{8})}\] is that the equation

OpenStudy (anonymous):

yes @sodukogirl13 execept only the sign doesnt go over all the numbers only the 2

OpenStudy (anonymous):

ok \[\sqrt{2}(5+\sqrt{8})\] \[\sqrt{10}+\sqrt{16}\] \[\sqrt{26}\]

OpenStudy (cggurumanjunath):